# UINO-FSS: Unifying Representation Learning and Few-shot Segmentation via Hierarchical Distillation and Mamba-HyperCorrelation

### v3

@@ I Introduction
Inspired by  [4], early works [5, 6] employ meta-learning to learn generalized matching models through extensive meta-tasks constructed from support-query pairs. Such a design, which closely mimics the few-shot inference, enables the learned model to be directly applied to segment novel classes without finetuning during inference. To enable effective support-query matching, prototype-based methods [7, 8, 9, 10, 11, 12] simply match the class prototype vector with each query pixel, while approaches based on embedding aggregation [13, 14, 15, 16, 17] extensively fuse the support and query feature maps for subsequent segmentation. Here, for embedding aggregation, convolutional hypercorrelation [16] and diverse attention-based matching [13, 14, 15, 18] are applied. Despite these efforts, FSS still suffers heavily from the inherent limitation of data scarcity and faces challenges in achieving stable and distinguished embeddings of novel classes.
@@
Particularly, existing FSS methods based on foundation models face the following challenges: 1) Hybrid frameworks that run two separate parallel branches would inevitable lead to high computational and memory costs, resulting in low efficiency and limited development in resource-constrained scenarios; 2) Beyond computation, SAM-based FSS methods still primarily rely on geometric prompts (e.g. points or boxes), especially in training-free solutions [24, 1, 2]. These methods could be sensitive to the point prompt selection. To overcome these issues, existing methods designs bi-directional prompt selection [1] or the graph-based segment filtering and merging processes [2], which, however, also introduce more hyper-parameters; 3) ViT self-attention lets each patch aggregate global context, yet co-occurring objects in the support image would also contaminate the support features due to the embedding coupling, where embeddings of one object incorporate semantics from other co-occurring objects.
@@ II-C Vision Foundation Models in Few-shot Segmentation
With the development of pre-trained visual models such as ViT [48], CLIP [21], SAM [22], and DINOv2 [20], Few-shot Semantic Segmentation (FSS) models [1, 49, 3] have achieved significant improvements on benchmark datasets. Particularly, [50] casts a CLIP on top of existing well-trained FSS networks, leading to further improvements. Among methods derived from SAM, the training-free method [1] generates point prompts using support-query similarity based on DINOv2 features, with dual-directional pruning. Inspired by [1][2] enhances segmentation by building a graph on the generated masks. Other works, like [3], aim to learn prompt encoders to avoid complex hand-crafted procedures but still rely on both SAM and DINOv2, resulting in large memory footprints. In contrast, [25] bypasses SAM and trains a decoder directly on DINOv2 features. Compared with [25], while both methods demonstrate high efficiency, but our method achieves significantly superior performance. This primarily stems from the pivotal finding of feature alignment between DINOv2's low-level features and SAM encoder's outputs, which enables a superior decoder through embedding distillation. Furthermore, our approach is enhanced by an effective MVPG module that generates both sparse prompts and Mamba-based dense prompts for decoding.
@@ III Preliminary
Since both DINOv2 and SAM are utilized for the few-shot segmentation task [1, 2, 3], the input images are passed through two separate image encoders to extract features independently. This results in a model with a large number of parameters and high inference time, making it less practical for real-world applications. To address this issue, we propose learning SAM's feature distribution by introducing a small number of trainable parameters after DINOv2 encoder, enabling SAM's powerful decoder on top of DINOv2. However, experimental results reveal significant distinctions between the features from these two models, as illustrated in Fig. 2. The distillation loss that aims to align the two kinds of features does not decrease. To this end, we compute self-similarities of features from DINOv2 and SAM's encoders separately for analysis.

### v4

@@ I Introduction
Inspired by  [4], early works [5, 6] employ meta-learning to learn generalized matching models through extensive meta-tasks constructed from support-query pairs. Such a design, which closely mimics the few-shot inference, enables the learned model to be directly applied to segment novel classes without finetuning during inference. To enable effective support-query matching, prototype-based methods [7, 8, 9, 10, 11, 12, 13] simply match the class prototype vector with each query pixel, while approaches based on embedding aggregation [14, 15, 16, 17, 18] extensively fuse the support and query feature maps for subsequent segmentation. Here, for embedding aggregation, convolutional hypercorrelation [17] and diverse attention-based matching [14, 15, 16, 19] are applied. Despite these efforts, FSS still suffers heavily from the inherent limitation of data scarcity and faces challenges in achieving stable and distinguished embeddings of novel classes.
@@
Particularly, existing FSS methods based on foundation models face the following challenges: 1) Hybrid frameworks that run two separate parallel branches would inevitably lead to high computational and memory costs, resulting in low efficiency and limited development in resource-constrained scenarios; 2) Beyond computation, SAM-based FSS methods still primarily rely on geometric prompts (e.g. points or boxes), which could be sensitive to the point prompt selection. To overcome these issues, [1] designs bi-directional prompt selection to prune noisy prompts. Later work [2] constructs a graph for segment filtering and merging processes, which, however, also introduces more hyper-parameters; 3) ViT self-attention lets each patch aggregate global context, yet co-occurring objects in the support image would also contaminate the support features due to the embedding coupling, where embeddings of one object incorporate semantics from other co-occurring objects.
@@ II-C Vision Foundation Models in Few-shot Segmentation
With the development of pre-trained visual models such as ViT [49], CLIP [22], SAM [23], and DINOv2 [21], Few-shot Semantic Segmentation (FSS) models [1, 50, 3] have achieved significant improvements on benchmark datasets. Particularly, [51] casts a CLIP on top of existing well-trained FSS networks, leading to further improvements. The training-free methods [25, 1] pioneer a SAM-based framework for FSS, where Matcher [1] generates point prompts using support-query similarity based on DINOv2 features, with dual-directional pruning. Inspired by [1][2] enhances segmentation by building a graph on the generated masks. Other works, like [3], aim to learn prompt encoders to avoid complex hand-crafted procedures but still rely on both SAM and DINOv2, resulting in large memory footprints. In contrast, [26] bypasses SAM and trains a decoder directly on DINOv2 features. Compared with [26], while both methods demonstrate high efficiency, but our method achieves significantly superior performance. This primarily stems from the pivotal finding of feature alignment between DINOv2's low-level features and SAM encoder's outputs, which enables a superior decoder through embedding distillation. Furthermore, our approach is enhanced by an effective MVPG module that generates both sparse prompts and Mamba-based dense prompts for decoding.
@@ III Preliminary
While existing methods [1, 2, 3] successfully harness the complementary strengths of SAM and DINOv2 by employing dual encoders for independent feature extraction, this design also leads to models with substantial parameter counts and high inference time. To address this issue, we propose learning SAM's feature distribution by introducing a small number of trainable parameters after DINOv2 encoder, enabling SAM's powerful decoder on top of DINOv2. However, experimental results reveal significant distinctions between the features from these two models, as illustrated in Fig. 2. The distillation loss that aims to align the two kinds of features does not decrease. To this end, we compute self-similarities of features from DINOv2 and SAM's encoders separately for analysis.